OpenStudy (anonymous):

Simplify 3x + 3 + 2. 3x + 5 5x + 3 8x

OpenStudy (anonymous):

combine the constants to get the answer

OpenStudy (anonymous):

i got 8x.

OpenStudy (anonymous):

The answer is A.

OpenStudy (anonymous):

It is A. 3 + 2 = 5. so then you have 3x + 5

OpenStudy (anonymous):

3x + 3 + 2 COMBINE LIKE TERMS: 3x + 5

OpenStudy (anonymous):

Add the common factors 3x+5

OpenStudy (mathstudent55):

You can't combine together numbers with x and numbers without x.

OpenStudy (mathstudent55):

You need to know what like terms are. Like terms are terms that have the same variable part. Only like terms can be combined together.

OpenStudy (anonymous):

thanks guys for the help!

OpenStudy (mathstudent55):

For example, 2x + 5 + 3x + 4 In this expression, 2x and 3x are like terms and can be combined together. 2x + 3x = 5x Also, 5 and 4 are like terms and can be combined together. 5 + 4 = 9 The final result is 5x + 9. 5x and 9 are not like terms because the variable parts of 5x and 9 are different. 5x has an x and 9 has no x. Since 5x and 9 are not like terms, they can't be combined together.

OpenStudy (mathstudent55):

wlcm